Nanotech, Inc., has a bond issue maturing in seven years that is paying a coupon rate of 10.51 percent (semiannual payments). Management wants to retire a portion of the issue by buying the securities in the open market. If it can refinance at 9.78 percent, how much will Nanotech pay to buy back its current outstanding bonds? (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and final answer to 2 decimal places, e.g. 15.25.)
Answer #1

Face Value of Bond = $1000

Semi-annual Coupon paymnet = $1000*10.51%*1/2

= $52.55

n = 7 yrs*2 =14

Market Interest Yield(YTM) = 9.78%

Semi-annual YTM = 9.78%/2 =4.89%

Calculating the Current Price of bond:-

+ ... Coupon Payment (1+YTM) Price Coupon Payment Coupon Payment + (1+YTM) (1 + YTM2 FaceValue (1+YTM)

Price = \frac{52.55}{(1+0.0489)^{1}}+\frac{52.55}{(1+0.0489)^{2}}+.....+\frac{52.55}{(1+0.0489)^{14}}+\frac{1000}{(1+0.0489)^{14}}

Price = $523.85 + $512.53

Price = $1036.38

So, Nanotech have to pay to buy back bonds at $1036.38 per bond

If you need any clarification, you can ask in comments.     

If you like my answer, then please up-vote as it will be motivating